OpenStudy (anonymous):

Find the zeros of the function: -12n^2 =-30n help please!

OpenStudy (anonymous):

n = 5/2 n = 0

OpenStudy (anonymous):

Can someone explain it please?

OpenStudy (anonymous):

@DaWizjr please explain?

OpenStudy (anonymous):

I cant

OpenStudy (anonymous):

@DaWizjr why not?

OpenStudy (anonymous):

@bahupepe help??

OpenStudy (anonymous):

Becuase

OpenStudy (anonymous):

ok wait so log here

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

-12n^2 +30n=0 f(n)= -12n(n-2/5) n=0 and n=2/5 means f (0) = 0 and  f (2/5) = 0

OpenStudy (anonymous):

Yea what she said

OpenStudy (anonymous):

Sane thing I said